OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

use point slope form

OpenStudy (anonymous):

y – y1 = m(x – x1) y+1=4(x+3)

OpenStudy (anonymous):

I just don't know how to do this crap -_-

OpenStudy (anonymous):

well you use point slope form it is just a formula you might wanna remember it if not google it but it will definately be on a test :P y – y1 = m(x – x1) since you are given a point and slope (-3, -1) with a slope of 4. you just plug in